Docstoc

SUS_Notes_10_12_14

Document Sample
SUS_Notes_10_12_14 Powered By Docstoc
					                                          Weekly SUS Meeting
                         Agenda & Summary Notes from 14 December, 2010
         Adv LIGO Suspension wiki: http://ilog.ligo-wa.caltech.edu:7285/advligo/Suspensions
                                                                 Every Tuesday at 9am Pacific/Noon Eastern/4pm UK
          MEETING AGENDA
1         LASTI/SEI/SUS schedule coordination discussion
2         Report from Technical/Fiber Meeting
3         Quad/BS/FM Mechanical - site build updates
4         Monolithic
5         OSEMs
6         Harnesses
7         Electronics


          Attendees
1         Norna, Eric                                                      Caltech
2         Russell, Ken, Allen, Liam, Marielle, Danny, Travis               Glasgow
4         Justin                                                           RAL
5         Brett, Rai                                                       MIT
7         Stuart                                                           Birmingham
8         Mike, Derek, Virginia, Bobby, Janeen                             LLO
          Mark, Betsy, Rob Lane, Andreas, Gerardo                          LHO
                                                                  SUMMARY NOTES AND ACTIONS FROM THIS MEETING
1 SUS/ISI LASTI/SEI/SUS Schedule Coordination Discussion
          Brett reports that the HAM ISI has arrived.

2         Report from Technical/Fiber Meeting                 http://ilog.ligo-wa.caltech.edu:7285/advligo/Suspensions/Technical_meetings
          Discussion on secondary prism placement. Rai presented G1001148-v1 and T0900388. He reocmmends using a phonograph cartridge to
          determine optimum position for secondary prism in both directions (vertical & front-to-back.) Russell notes that the design they proposed at the review was to
          glue the primary and secondary prisms at the same time. Calum suggests adding to the procedure the testing of the secondary prism position. Rai said it is not
          clear how sensitive the positioning is; we need to determine how broad this range is.
3         Mechanical:

          LHO Site Build status - Betsy reports of the weird pitch mode drop from .5 Hz to .41 Hz, which leapfrogs the pendulum mode.
          More mass was added to the test mass and penultimate mass. Norna asks Joe to analyze in ProE whether the center of mass has moved.


          LLO Site Build status - The second quad is undergoing testing now. It has the same pitch mystery as the LHO quad.


4         Monolithic Stage:                       http://ilog.ligo-wa.caltech.edu:7285/advligo/Monolithic_Suspensions/Meetings
          Travis & Danny are at Glasgow now for welding training.
          Marielle has updated the prism bonding procedure, E1000752 for the ERM & TCP. She would like Betsy & Margot to review and send comments.
          Doug reports that they're working on two bugs in the fiber pulling. Just an issue with hardware wiring. Jessie is working this issue. Bluetooth wireless
          keyboard seems to work with everything but Labview. They pulled two fibers and showed to NSF.

5         Osems
          BOSEMs - Stuart reports he's received 20 BOSEMs for retrofitting. They've accomplished about 80 retrofits at this point.


          AOSEMs - Jay reports that 360 AOSEMs have been sent to LHO for processing so far. Betsy reports that 143 had filters glued and she has another
          100 cleaned and ready to have their filters glued. Jay reports that there are another 103 coming to LHO; they're awaiting flexi-circuits.
          Noise calculations have been done by Mark and Jay and they meet the requirements.
6         Electronics -
          Ron reports: Transmons: all complete,OMCs: all complete. PUMs: The last 4 PUM boards should have all their 797s replaced by AD8671s by the end of today.
          All 12 satellite boxes required for the PUMs have now been tested. 2 spare monitor boards are in test.
          3 PUM Drive boards tested. 1 more is in test. 3 units (numbers 5, 6, and 7) are being built.



7         Harnesses -




          Agenda for 21 December, 2010
          1.   LASTI/SEI/SUS schedule coordination
          2.   Report from Technical/Fiber Meeting
          3.   Quad/BS/FM Mechanical and site build update
          4.   Monolithic
          5.   OSEMs
          6.   Electronics
          7.   Harnesses

				
DOCUMENT INFO
Shared By:
Categories:
Tags:
Stats:
views:0
posted:12/21/2011
language:
pages:1